OpenStudy (anonymous):

5. Solve the equation. Show your work. Check your answer. 4y + 5 = – 31 Here's what i go i'm not sure if i'm wrong or right 4y + 5 = 31 -4 -4 +5 = -35 5/5 -35/5 35/5 = 7

TheSmartOne (thesmartone):

Wrong.

TheSmartOne (thesmartone):

\(\sf\Large 4y + 5 = – 31\) First we need to take 5 to the other side. So we need to subtract it on both sides... \(\sf\Large4y + 5 -5 = -31 -5\)

OpenStudy (anonymous):

y=-36/4=-9

OpenStudy (anonymous):

Sorry was away from key >.<

OpenStudy (anonymous):

Okay thanks Both of ya'll. Since TheSmartOne explained the answer i'll give them the medal.

TheSmartOne (thesmartone):

:)